OpenStudy (anonymous):

What is (3xy-2b)(3xy+2b)

OpenStudy (alexandervonhumboldt2):

hint: (a-b)*(a+b)=a^2-b^2

OpenStudy (alexandervonhumboldt2):

a=3xy b=2b

OpenStudy (anonymous):

9xy-4b??? Totally illiterate by the way....😰😂

OpenStudy (mathstudent55):

You multiplied the numbers correctly, and you got the sign correct. You also need to multiply the variables. \((3xy-2b)(3xy+2b)\) \(= (3xy)^2 - (2b)^2\) \(= 3xy \times 3xy - 2b \times 2b\) \(= 3^2 x^2 y^2 - 2^2 b^2\) \(=9x^2y^2 - 4b^2\)

OpenStudy (anonymous):

Aww- thank you!!

OpenStudy (mathstudent55):

You're welcome.